The isotope H-3 has a half life of 7.02 years. If you start with an initial...

  1. The isotope H-3 has a half life of 7.02 years. If you start with an initial H-3 mass of 100.0g. What is the approximate amount remaining after 49 years of decay?
Answer #1

Answer=

The given are

The isotope H-3 has a half life (t1/2) =  7.02 years.

T = 49 years

If you start with an initial H-3 mass of 100.0g.

Initial reactivity N0 = 100 gm

T = n X t1/2

n = T / t1/2

n = 49 / 7.02

= 6.980

Formula is

N = N0 (1/2)n

= 100 ( 1/2 ) 6.98

= 100 X (1/126.23 )

=0.792

The approximate amount remaining after 49 years of decay = 0.792 gms .